OpenStudy (anonymous):

Solve for x: logx+log(x-1)=log(4x) i got x=5 ... is this correct?

OpenStudy (jennychan12):

x^2-x = 4x x^2-5x=0 x(x-5) = 0 it's 0 or 5 but since you can't have a negative inside the log, answer is five. correct :D

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

here log(x(x-1))=log4x or x(x-1)=4x or x^2-x=4x or x^2 -5x=0 hence x(x-5)=0 hence x=0 or x=5 but if x=0 we have log(x-1)=log(-1) (absurd as we cannot take log of negatives) hence x=5